Determine the Laplace transform of the given generalized function. 58(t - 4) L{58(t - 4)}(s) =
Answer #1

L{si(4-1)} (x) = 5 £{8 (t-4)}() (: [{S(-e)]=ées) L{58 (1-4)} (s) = seus =5 e ė45